OpenStudy (anonymous):

27. Christina takes photos for $100 an hour. She worked 8.5 hours on Monday. She deposited half of what she earned into her savings account. She worked another 5 hours on Wednesday. She stopped and filled up her car with $51.70 worth of gas and deposited the rest. If she withdraws $85.25, how much is left in her savings account?

OpenStudy (anonymous):

100 dollars an hour. Monday she worked 8.5 hours 100*8.5= 850 dollars 850/2= 425 dollars Wednesday she worked 5 hours. 100*5= 500 dollars 500-51.70= 448.30 Deposits 448.30 Savings Account is now at 873.30 Withdraws 85.25 873.30-85.25= 788.05 Answer is 788.05